The sum of two numbers is 9. The greater number is 3 more than twice the lesser number. What are the numbers?
Answer by rm29924(97)

let x = the lesser number
the greater number is 3 more than twice the lesser number
twice the lesser number is 2x
3 more would be 2x+3
the greater number plus the lesser number = 9
x+2x+3=9
combine like terms
3x+3=9
subtract 3 from both sides
3x=6
divide both sides by 3
x=2
answer: the lesser number is 2 and the greater number is 2x+3 2(2)+3
4+3=7 so the greater numer is 7
check answer: 2+2x+3=9
2+2(2)+3=9
2+4+3=9
9=9